A marathon race is 26 miles, 385 yards. What is this distance in rods, furlongs, meters, and kilometers?
Answer #1

It just a unit conversion
(a)

1 mile = 320 rods

26 miles = 26 x 320 = 8320 rods

1 yard = 0.181818 rods

385 yards = 70 rods

total distance = 8320 + 70 = 8390 rods

(b)

1 mile = 8 us furlongs

26 miles = 8 x 26 = 208 furlongs

1 yard = 0.00454545 furlongs

385 yards = 1.75 furlongs

total distance = 208 + 1.75 = 209.75 furlongs

(c)

1 mile = 1609.34m

26 miles = 41842.9m

1 yard = 0.9144 meters

385 yards =  meters

total distance = 352.044m

total distance = 352.044m + 41842.9m = 42194.94 m

(d)

1000m = 1km

42194.94 m = 26 miles = 42.195 km